On planet X, an object falls 18 feet in 3 seconds. Knowing the distance it falls varies directly with the square of the time of fall, how long does it take an object to fall 100 feet? Round your answer to three decimal places
Answer by stanbon(75887)

On planet X, an object falls 18 feet in 3 seconds. Knowing the distance it falls varies directly with the square of the time of fall, how long does it take an object to fall 100 feet? Round your answer to three decimal places
---------------
distance = k*time^2
d = kt^2
Find "k":
18 = k*3^2
k = 2
----------
Equation for planet X:
d = 2t^2
100 = 2*t^2
t^2 = 50
t = 5sqrt(2)
t = 7.071 seconds
=====================
Cheers,
Stan H.